dS dt
=1000r 100 e(r/100)
The rate of change of the value of an investment, S, with respect to time, t ≥ 0, is given by
where r is the annual interest rate (assumed constant) and the principal of the investment is S(0) = 1 000
1.(a) Find an expression for S(t), that is, the value of the investment at time t.
1
Expert's answer
2020-06-10T18:50:27-0400

dS/dt = 1000("\frac {r} {100}" )"e^{r\/100}"

dS = 10r"e^{r\/100}" dt

Integrating with respect to t

"\int" dS = "\int" 10r"e^{r\/100}" dt +C , where C is integration constant.

S = 10r"e^{r\/100}" t + C

When t = 0, S =1000

So 1000 = C

Therefore S = 10r"e^{r\/100}" t + 1000

The value of investment at time t is

S(t) = 10r"e^{r\/100}" t + 1000


"\mathbb { NOTE}"

As per growth rate formula if a correction is made as

"\frac {dS}{dt}" = 1000("\frac {r}{100}" ) "e^{rt\/100}"

"\int" dS = "\int" 10r"e^{rt\/100}" dt

=> S = 10r"e^{rt\/100}" ."\frac {100}{r}" + C

=> S = 1000"e^{rt\/100}" +C

When t= 0 , S = 1000

=> 1000 = 1000 + C

=> C = 0

So S = 1000"e^{rt\/100}"

The value of investment at time t is

S(t) = 1000"e^{rt\/100}"